OpenStudy (anonymous):

-5n (-9n + 1) help Anyone???

jimthompson5910 (jim_thompson5910):

-5n times -9n is ???

OpenStudy (anonymous):

45n

jimthompson5910 (jim_thompson5910):

n*n = n^2

jimthompson5910 (jim_thompson5910):

so it's really 45n^2

jimthompson5910 (jim_thompson5910):

-5n times 1 = ???

OpenStudy (anonymous):

-5n?

jimthompson5910 (jim_thompson5910):

so -5n (-9n + 1) -5n (-9n ) - 5n(1) -5n times -9n - 5n times 1 45n^2 - 5n So \[\Large -5n (-9n + 1) = 45n^2 - 5n\]

OpenStudy (anonymous):

thank you so much

jimthompson5910 (jim_thompson5910):

sure thing